OBSERVATION AND ANALYSIS OF NEW HIGH-J INTER-SPECIES TRANSITIONS IN CH_{2}DOH

\underline{INDRANATH MUKHOPADHYAY}, DAVID. S PERRY, KNIGHT CHEMICAL LABORATORY, UNIVERSITY OF AKRON, AKRON, OH, 44325; JOHN C. PEARSON, JET PROPULSION LAB-ORATORY, CALIFORNIA INSTITUTE OF TECHNOLOGY, PASADENA, CA 91109-8099; YUN-BO DUAN, DEPARTMENT OF PHYSICS, UNIVERSITY OF BRITISH COLUMBIA, VANCOU-VER, BC, CANADA; REBECCA A.H. BUTLER, ERIC HERBST, AND FRANK C. DE LUCIA, DEPARTMENT OF PHYSICS, THE OHIO STATE UNIVERSITY, COLUMBUS, OH 43210-1106.

We report the observation and assignments of several series of transitions in CH_{2}DOH, including a new and intense series of Q-branch c-type transitions between the e_{1} and o_{1} torsional sub-states through high values of the rotational quantum number J. Other transitions assigned are c-type R-branch and P -branch transitions between these two sub-states and a-type lines within the e_{1} and o_{1} sub-states. The assignments were facilitated by initial analysis of the strong Q-branch series, as recorded by the FAst Scan Sub-millimeter Spectroscopy Technique (FASSST). The assigned lines of CH_{2}DOH include the first in the sub-millimeter-wave and the first to possess high J-values. The completeness of the data generated by FASSST and the success of a simple power series analysis suggest that many parts of the spectrum of CH_{2}DOH may be far more tractable than previously believed. The data should be useful in the development of a full Hamiltonian and in the assignment of astronomical emission features.